# Infinitely many genus two and three curves of small fixed positive rank

### v2

@@ 2.2. Partial rational two-torsion
Now we consider the family of bielliptic curves $C_{m,d}\colon y^{2}=d^{3}x^{6}+m^{3}$, where $d,m\in\mathbb{Z}$ are squarefree. Similarly as before, [BD11, Theorem 3.2] implies that $\mathrm{Jac}(C_{m,d})$ is isogenous to $E_{d}\times E_{m}$, where $E_{d}\colon y^{2}=x^{3}+d^{3}$ and $E_{m}\colon y^{2}=x^{3}+m^{3}$, which is a quadratic twist of $E_{d}$ by $\frac{m}{d}$. Now, we fix $d$ and vary $m$, obtaining the following theorem.
@@
We note that this approach works only over $\mathbb{Q}$, but unlike the previous approach, it gives explicit families of infinitely many non-isomorphic curves of fixed positive rank. This family is an infinite set because the curve $C_{p,d}$ has a place of bad reduction at $p$. A quick Magma search gives us the following rank information:
@@ 2.3. Full rational two-torsion - over $\mathbb{Q}$
which after replacing $(x,y)$ by $(px,p^{2}y)$ becomes $E_{2,d,k}\colon y^{2}=d(x-k)(x^{3}-x)$. From now on, we can search directly in Magma for curves $E_{2,d,k}$ of any rank we want. Still, if we prefer to use the standard way of expressing elliptic curves, a linear change of variables leads to the following model
@@
Let $r\in\mathbb{Z}_{\geq 0}$, $k\in\mathbb{Z}\setminus\{-1,0,1\}$, and squarefree $d\in\mathbb{Z}$ be such that $\mathrm{Rank}(E^{\prime}_{2,d,k}(\mathbb{Q}))=r$, where $E^{\prime}_{2,d,k}\colon dy^{2}=(x+k^{2}-k)(x+k^{2}-1)(x+k^{2}+k)$. Then, for the following curves of genus two
@@
where $p\equiv 3\pmod{8}$ is a prime number, we have $\mathrm{Rank}(C_{d,k,p}(\mathbb{Q}))=r$.
We note that given a fixed value of $k$ and $d$, the curve $C_{d,k,p}$ has a place of bad reduction at $p$. Hence, the collection of curves $\{C_{d,k,p}\}_{p\equiv 3\mod 8}$ is an infinite set. Here also we obtain results over $\mathbb{Q}$, but we can give some interesting explicit examples. Testing the curves $E^{\prime}_{2,d,k}$ for $d=1$ in Magma, we obtain the following rank information:
@@ 2.4. Full rational two-torsion - over any number field
Combining the above theorem with recent breakthroughs in distribution of algebraic ranks of quadratic twist families of elliptic curves over number fields [Smi22b, KP25], we can prove the following result for genus 2 curves over any number field $K$.
@@
  • #$\#\mathcal{O}_{K}/(p_{i}),\#\mathcal{O}_{K}/(q_{i}),\#\mathcal{O}_{K}/(r_{i})>5$.
@@
Now we apply Lemma 2.9 to the ideal $I=(8)$ and $\rho=1$ to find totally positive prime elements $p_{1},p_{2},p_{3}\in\mathcal{O}_{K}$ such that #$\#\mathcal{O}_{K}/(p_{i})>5$. Then, we apply Lemma 2.9 to the ideal $I=(8p_{1}p_{2}p_{3})$ and $\rho=1$ to find totally positive prime elements $q_{1},q_{2},q_{3}\in\mathcal{O}_{K}$ such that #$\#\mathcal{O}_{K}/(q_{i})>5$. Finally, we apply Lemma 2.9 to the ideal $I=(8p_{1}p_{2}p_{3}q_{1}q_{2}q_{3})$ and $\rho=1$ to find totally positive prime elements $r_{1},r_{2},r_{3}\in\mathcal{O}_{K}$ such that #$\#\mathcal{O}_{K}/(r_{i})>5$. By construction, we have that $\left(\dfrac{p_{i}}{q_{j}}\right)=\left(\dfrac{q_{j}}{r_{k}}\right)=\left(\dfrac{r_{k}}{p_{i}}\right)=1$.
@@
By [Smi22b, Theorem 2.14, Example 3.1], there exists infinitely many (in fact a positive proportion of) square-free elements $m,d\in\mathcal{O}_{K}$ such that $\mathrm{Rank}(E_{1,m}(K))=\mathrm{Rank}(E_{2,d}(K))=0$. For such choices of $m,d\in\mathcal{O}_{K}$ we have $\mathrm{Rank}(\mathrm{Jac}(C_{m,d})(K))=0$.
We turn our focus to constructing infinitely many curves $C_{m,d}$ such that $\mathrm{Rank}(\mathrm{Jac}(C_{m,d})(K))=1$. By our choice of nine prime elements $p_{i},q_{j},r_{k}$ the elliptic curve $E_{1,1}$ satisfies the following three conditions:
@@
3. Genus three curves
Similar to the genus two case, we consider bielliptic curve $C_{3}$ of genus $3$ Its Jacobian variety decomposes as $\mathrm{Jac}(C_{3})\sim E\times\mathrm{Jac}(C_{2})$, where $C_{2}$ is a genus two curve over $K$. Instead of working over general number fields, we restrict our focus to $K=\mathbb{Q}$. In a similar spirit to Theorem 2.1, we show that as long as there exists a curve $C_{2}$ with full rational 2-torsion whose Jacobian has algebraic rank $r$ one can vary $E$ to construct infinitely many curves of genus three and rank $r$.
@@ 3. Genus three curves
Let $r\in\mathbb{Z}_{\geq 0}$. Let $a,b,c,d\in\mathbb{Q}^{\times}$ be distinct rational numbers. Assume that there is a hyperelliptic curve $C_{2}:y^{2}=x(x-a)(x-b)(x-c)(x-d)$ such that $\mathrm{Rank}(\mathrm{Jac}(C)(\mathbb{Q}))=r$. Then there are infinitely many genus three curves $C_{3,D}:Dy^{2}=(x^{2}-Da)(x^{2}-Db)(x^{2}-Dc)(x^{2}-Dd)$ defined over $\mathbb{Q}$ such that $\mathrm{Rank}(\mathrm{Jac}(C_{3,D})(\mathbb{Q}))=r$.
@@
Given any $a^{\prime},b^{\prime},c^{\prime},d^{\prime}\in\mathbb{Q}^{\times}$ and a square-free integer $D\in\mathbb{Z}$, consider the following three curves over $\mathbb{Q}$.
$\begin{split}C_{3,D}&:Dy^{2}=(x^{2}-a^{\prime})(x^{2}-b^{\prime})(x^{2}-c^{\prime})(x^{2}-d^{\prime})\\ E_{D}&:Dy^{2}=(x+(d^{\prime}-a^{\prime})(d^{\prime}-b^{\prime}))(x+(d^{\prime}-a^{\prime})(d^{\prime}-c^{\prime}))(x+(d^{\prime}-b^{\prime})(d^{\prime}-c^{\prime}))\\ C_{2,D}&:Dy^{2}=x(x-a^{\prime})(x-b^{\prime})(x-c^{\prime})(x-d^{\prime}).\end{split}$ (4)
Then there exists an isogeny $\phi:\mathrm{Jac}(C_{3,D})\to E_{D}\times\mathrm{Jac}(C_{2,D})$, which follows directly from [Sha15, Theorem 1]. We note that the theorem is proven for genus three curves over $\mathbb{C}$, which can be generalized to genus three curves over $\mathbb{Q}$ we consider in this proof, see the end of [Sha15, Introduction]. The change of variables one needs to choose can be given as follows. For the elliptic curve $E_{D}$, we use the change of variables $\tilde{x}=x^{2}$ and $\tilde{y}=y$ to obtain Weierstrass equation for $E^{\prime}_{D}:D\tilde{y}^{2}=(\tilde{x}-a^{\prime})(\tilde{x}-b^{\prime})(\tilde{x}-c^{\prime})(\tilde{x}-d^{\prime})$. Afterwards, we use $X:=\frac{1}{(\tilde{x}-d^{\prime})(d^{\prime}-a^{\prime})(d^{\prime}-b^{\prime})(d^{\prime}-c^{\prime})}$ and $Y:=\tilde{y}\tilde{x}^{2}(d^{\prime}-a^{\prime})^{2}(d^{\prime}-b^{\prime})^{2}(d^{\prime}-c^{\prime})^{2}$ to obtain the Weierstrass equation for $E_{D}$. For the hyperelliptic curve $C_{2,D}$ we use the change of variables $X:=x^{2}$ and $Y:=xy$. We hence obtain that the function field of $C_{3,D}$, which is $\mathbb{Q}(x,y)$, contains a subfield $\mathbb{Q}(x^{2},y)$ corresponding to $E_{D}$ and another subfield $\mathbb{Q}(x^{2},xy)$ corresponding to $C_{2,D}$.
We now construct our desired infinite family of genus 3 curves. Let $a,b,c,d$ be integers which satisfy the following equation:
[…]
With this choice of parameters, the Weierstrass equations for the three curves can be rewritten as:
$\begin{split}C_{3,D}&:Dy^{2}=(x^{2}-Da)(x^{2}-Db)(x^{2}-Dc)(x^{2}-Dd),\\ E_{D}&:Dy^{2}=(x+(d-a)(d-b))(x+(d-a)(d-c))(x+(d-b)(d-c)),\\ C_{2}&:y^{2}=x(x-a)(x-b)(x-c)(x-d).\end{split}$ (6)
We note that the collection $\{C_{3,D}\}_{D\text{ square-free}}$ contains an infinite number of non-isomorphic genus $3$ curves over $\mathbb{Q}$, because as long as $D$ is coprime to $(d-a)(d-b)(d-c)(a-b)(b-c)(c-a)$, the elliptic curve $E_{D}$ has additive reduction over every prime divisors of $D$. By [Smi25, Theorem 1.1], there are infinitely many (in fact half of the) square-free integers $D$ such that $E_{D}$ satisfies $\mathrm{Rank}(E(\mathbb{Q}))=0$. (In fact, if one can guarantee that none of the triples of integers $(d-a)(d-b)(c-b)(c-a)$, $(d-a)(d-c)(b-c)(b-a)$, and $(d-b)(d-c)(a-c)(a-b)$ are squares, then one can use [Kan13] to show that a positive proportion of squarefree integers $D$ satisfies $\mathrm{Rank}(E_{D}(\mathbb{Q}))=0$). This implies that for infinitely many such choices of $D$, we have $\mathrm{Rank}(\mathrm{Jac}(C_{3,D})(\mathbb{Q}))=r$. ∎
As a consequence, we prove that there are infinitely many genus three hyperelliptic curves defined over $\mathbb{Q}$, whose rank of the Jacobian over $\mathbb{Q}$ is a fixed number between 0 and 4. We are especially thankful to Jennifer Balakrishnan for providing us with examples of genus two curves with full rational two torsion, whose Jacobians have algebraic rank equal to 3 and 4.
@@
Let $0\leq r\leq 4$ be a fixed integer. Then there are infinitely many genus three curves $C$ defined over $\mathbb{Q}$ such that $\mathrm{Rank}(\mathrm{Jac}(C)(\mathbb{Q}))=r$.

### v3

@@ 2.2. Partial rational two-torsion
Now we consider the family of bielliptic curves $C_{m,d}\colon y^{2}=d^{3}x^{6}+m^{3}$, where $d,m\in\mathbb{Z}$ are square-free. Similarly as before, [BD11, Theorem 3.2] implies that $\mathrm{Jac}(C_{m,d})$ is isogenous to $E_{d}\times E_{m}$, where $E_{d}\colon y^{2}=x^{3}+d^{3}$ and $E_{m}\colon y^{2}=x^{3}+m^{3}$, which is a quadratic twist of $E_{d}$ by $\frac{m}{d}$. Now, we fix $d$ and vary $m$, obtaining the following theorem.
@@
We note that this approach works only over $\mathbb{Q}$, but unlike the previous approach, it gives explicit families of infinitely many curves of fixed positive rank. A quick Magma search gives us the following rank information:
@@ 2.3. Full rational two-torsion - over $\mathbb{Q}$
which after replacing $(x,y)$ by $(px,p^{2}y)$ becomes $E_{2,d,k}\colon y^{2}=d(x-k)(x^{3}-x)$. From now on, we can search directly in Magma for curves $E_{2,d,k}$ of any rank we want. Still, if we prefer to use the standard way of expressing elliptic curves, a rational change of variables leads to the following model
@@
Let $r\in\mathbb{Z}_{\geq 0}$, $k\in\mathbb{Z}\setminus\{-1,0,1\}$, and square-free $d\in\mathbb{Z}$ be such that $\mathrm{Rank}(E^{\prime}_{2,d,k}(\mathbb{Q}))=r$, where $E^{\prime}_{2,d,k}\colon dy^{2}=(x+k^{2}-k)(x+k^{2}-1)(x+k^{2}+k)$. Then, for the following curves of genus two
@@
where $p\equiv 3\pmod{8}$ is a prime number, we have $\mathrm{Rank}(\mathrm{Jac}(C_{d,k,p})(\mathbb{Q}))=r$.
Testing the curves $E^{\prime}_{2,d,k}$ for $d=1$ in Magma, we obtain the following rank information:
@@
2.4. Rational points on these curves
[…]
On the other hand, also by [Sil09, Exercise 10.19], for $E_{p}\colon y^{2}=x^{3}+p^{3}$ with $p\equiv 5\pmod{12}$, we have $E_{p}(\mathbb{Q})=\{O,(-p,0)\}$. Hence, if $d$ in a square (i.e. $d=1$ since we assume $d$ is square-free), then $C_{p,d}(\mathbb{Q})=\{\infty_{\pm}\}$. Otherwise, the set $C_{p,d}(\mathbb{Q})$ is empty. When $p\equiv 3\pmod{4}$, the problem of determining $C_{p,d}(\mathbb{Q})$ could be very difficult.
@@ 2.5. Full rational two-torsion - over any number field
Combining the above theorem with recent breakthroughs in distribution of algebraic rank of quadratic twist families of elliptic curves over number fields [Smi22b, KP25], we can prove the following result for genus 2 curves over any number field $K$.
@@
  • $\mathcal{O}_{K}/(p_{i}),\#\mathcal{O}_{K}/(q_{i}),\#\mathcal{O}_{K}/(r_{i})>5$.
@@
Now we apply Lemma 2.9 to the ideal $I=(8)$ and $\rho=1$ to find totally positive prime elements $p_{1},p_{2},p_{3}\in\mathcal{O}_{K}$ such that $\mathcal{O}_{K}/(p_{i})>5$. Then, we apply Lemma 2.9 to the ideal $I=(8p_{1}p_{2}p_{3})$ and $\rho=1$ to find totally positive prime elements $q_{1},q_{2},q_{3}\in\mathcal{O}_{K}$ such that $\mathcal{O}_{K}/(q_{i})>5$. Finally, we apply Lemma 2.9 to the ideal $I=(8p_{1}p_{2}p_{3}q_{1}q_{2}q_{3})$ and $\rho=1$ to find totally positive prime elements $r_{1},r_{2},r_{3}\in\mathcal{O}_{K}$ such that $\mathcal{O}_{K}/(r_{i})>5$. By construction, we have that $\left(\dfrac{p_{i}}{q_{j}}\right)=\left(\dfrac{q_{j}}{r_{k}}\right)=\left(\dfrac{r_{k}}{p_{i}}\right)=1$.
@@
By [Smi22b, Theorem 2.14, Example 3.1], there exists infinitely many (in fact a positive proportion of) square-free elements $m,d\in\mathcal{O}_{K}$ such that $\mathrm{Rank}(E_{1,m}(K))=\mathrm{Rank}(E_{2,d}(K))=0$. Here, we note that there is a subtlety with generating infinitely many such rank 0 quadratic twist families of elliptic curves over any number field $K$. This is where the “parity-invariant subcase” appearing after [Smi22b, Case 2.13] becomes relevant. (We would like to sincerely thank Adam Morgan for pointing out this subtlety.) Our construction of elliptic curves $E_{1,m}$ and $E_{2,d}$ shows that both curves have at least one place of multiplicative reduction. By the first paragraph of proof of [MS, Proposition 4.3] and [MS, Lemma 4.2], performing an unramified quadratic twist at the local condition over such a multiplicative place changes the parity of 2-Selmer group of $E_{1,m}$ (and likewise for $E_{2,d}$). Hence, we can always find infinitely many $m,d\in\mathcal{O}_{K}$ (in fact a positive density of square-free $m,d\in\mathcal{O}_{K}$) such that the parity of 2-Selmer group of $E_{1,m}$ is even. By [Smi22b, Theorem 2.14, Example 3.1], there are infinitely many choices of $m,d\in\mathcal{O}_{K}$ such that $\mathrm{Rank}(\mathrm{Jac}(C_{m,d})(K))=0$.
We turn our focus to constructing infinitely many curves $C_{m,d}$ such that $\mathrm{Rank}(\mathrm{Jac}(C_{m,d})(K))=1$. We note that this result for $g=2$ was obtained before [KM]. By our choice of nine prime elements $p_{i},q_{j},r_{k}$ the elliptic curve $E_{1,1}$ satisfies the following three conditions:
@@
3. Higher genus curves
We generalise the method of gluing two elliptic curves explored in Section 2 to gluing two hyperelliptic curves. This allows us to construct infinitely many hyperelliptic curves of any genus over any number field whose Jacobians have fixed positive rank between 0 and $2$. We also explore examples of genus 3, 4, 5, and 6 hyperelliptic curves over $\mathbb{Q}$ with fixed small positive rank.
3.1. Higher genus - over any number field
Let $K$ be a number field, $f=x^{n}+a_{n-1}x^{n-1}+\cdots+a_{1}x+a_{0}\in\mathcal{O}_{K}[x]$ a monic polynomial of degree $n$, and $d\in\mathcal{O}_{K}$. Define
$f_{d}(x)=x^{n}+da_{n-1}x^{n-1}+\cdots+d^{n-1}a_{1}x+d^{n}a_{0},$
i.e., the polynomial such that $f_{d}(dx)=d^{n}f(x)$ Consider the following three hyperelliptic curves
$C$ $\colon dy^{2}=f_{d}(x^{2});$
$C_{1}$ $\colon dy^{2}=xf_{d}(x);$
$C_{2}$ $\colon dy^{2}=f_{d}(x).$
The curve $C$ admits two degree 2 morphisms
$\psi_{1}\colon C\longrightarrow C_{1},$ $\;\;\psi_{1}(x,y)=(x^{2},xy);$
$\psi_{2}\colon C\longrightarrow C_{2},$ $\;\;\psi_{2}(x,y)=(x^{2},y).$
Lemma 3.1.
The map $\Psi=({\psi_{1}}_{*},{\psi_{2}}_{*})\colon\mathrm{Jac}(C)\longrightarrow\mathrm{Jac}(C_{1})\times\mathrm{Jac}(C_{2})$ is an isogeny defined over $K$.
Proof.
The map $\Psi$ is defined over $K$ as $\psi_{1},\psi_{2}$ are defined over $K$. We denote by $C^{(g(C))}$, $C_{1}^{(g(C_{1}))}$, and $C_{2}^{(g(C_{2}))}$ the symmetric powers of $C$, $C_{1}$, and $C_{2}$ by respective genus g(C)=n1g(C)=n-1, $g(C_{1})=\lfloor\frac{n}{2}\rfloor$ and $g(C_{2})=\lfloor\frac{n-1}{2}\rfloor$. Using [Mil86b, Theorem 5.1, Corollary 7.5], we have the following commutative diagram.
$C^{g}$$C_{1}^{g(C_{1})}\times C_{2}^{g(C_{2})}$$C^{(g(C))}$$C_{1}^{(g(C_{1}))}\times C_{2}^{(g(C_{2}))}$$\mathrm{Jac}(C)$$\mathrm{Jac}(C_{1})\times\mathrm{Jac}(C_{2})$$\scriptstyle{\psi_{1}\times\psi_{2}}$$\scriptstyle{\Psi}$
The surjectivity of $\Psi$ follows from the surjectivity of $C^{g}\rightarrow C_{1}^{g(C_{1})}\times C_{2}^{g(C_{2})}$ and all vertical maps in the commutative diagram. Because $\Psi$ is a homomorphism of abelian varieties and since $\lfloor\frac{n}{2}\rfloor+\lfloor\frac{n-1}{2}\rfloor=n-1$, we have $\dim\mathrm{Jac}(C)=\dim(\mathrm{Jac}(C_{1})\times\mathrm{Jac}(C_{2}))$, so, we can use [Mil86a, Proposition 8.1] to show that $\Psi$ is an isogeny. ∎
Remark 3.2.
Thanks to Maarten Derickx’s suggestion, we can note that the pull-backs $\psi_{1}^{*}(\mathrm{Jac}(C_{1}))\subseteq\mathrm{Jac}(C)$ and $\psi_{2}^{*}(\mathrm{Jac}(C_{2}))\subseteq\mathrm{Jac}(C)$ have 0-dimensional intersection inside $\mathrm{Jac}(C)$. This can be checked by pulling-back the holomorphic differentials $\frac{dx}{y},\ldots\frac{x^{\lfloor\frac{n}{2}\rfloor-1}dx}{y}$ from $\mathrm{Jac}(C_{1})$ and $\frac{dx}{y},\ldots\frac{x^{\lfloor\frac{n-1}{2}\rfloor-1}dx}{y}$ from $\mathrm{Jac}(C_{2})$, which give the full basis of the holomorphic differentials $\frac{dx}{y},\ldots\frac{x^{n-2}dx}{y}$ on $\mathrm{Jac}(C)$.
Note that, depending on the parity of $n$, after a linear change of variables ($x\mapsto dx$), we can simplify the equations for $C_{1}$ and $C_{2}$
$C_{1}\cong y^{2}=xf(x),\;C_{2}\cong dy^{2}=f(x)$ $\text{ if }2\mid n,$
$C_{1}\cong dy^{2}=xf(x),\;C_{2}\cong y^{2}=f(x)$ $\text{ if }2\nmid n.$
Note that in both cases, one of the curves is isomorphic to a fixed one, given by an odd degree model, whereas an even degree model becomes a twist by $d$. Hence, this gives a clear strategy to construct infinitely many curves of genus $g$ of a fixed rank $r$; we just need to find a specific one satisfying some mild conditions.
Before we proceed, we recall a few definitions and results from previous literature.
Definition 3.3 (Definition 1.2 and 2.2 of [KM]).
Let $C:y^{2}=(x-a_{1})\cdots(x-a_{2g+1})$ be an odd degree hyperelliptic curve over a number field $K$. Given a place $\omega$ of $K$ and a pair of distinct indices $1\leq i,j\leq 2g+1$, we say that $\omega$ is multiplicative of type $\{i,j\}$ if for every pair of distinct indices $1\leq k,\ell\leq 2g+1$, we have
(4) $\mathrm{val}_{\omega}(a_{k}-a_{\ell})=\begin{cases}1&\text{ if }\{k,\ell\}=\{i,j\},\\ 0&\text{ otherwise}.\end{cases}$
Given a positive number $B>0$, we say that $C$ is $B$-generic if there exist $4g^{2}+4g-1$ distinct finite places
$\begin{cases}\omega_{1},\omega_{2},&\cdots,\omega_{4g^{2}+2g-1},\\ \omega^{\prime}_{1},\omega^{\prime}_{2},&\cdots,\omega^{\prime}_{2g}.\end{cases}$
which satisfy the following four conditions.
  • For every $1\leq i,j\leq 2g$, the place $\omega_{(i-1)2g+j}$ is multiplicative of type $\{j,2g+1\}$.
  • For every $1\leq i\leq 2g-1$, the place $\omega_{4g^{2}+i}$ is multiplicative of type $\{i,i+1\}$.
  • For every $1\leq i\leq 4g^{2}+2g-1$, we have $\mathcal{O}_{K}/(\omega_{i})\geq B$.
  • For every $1\leq i\leq 2g$, the place $\omega^{\prime}_{i}$ is of type $\{i,2g+1\}$.
Definition 3.4.
Using the notation of Definition 3.3, we say that $C$ is BB-über-generic if the following conditions are satisfied: There are $g^{2}(2g+1)$ distinct places of $K$ such that for each pair $1\leq k<\ell\leq 2g+1$ we have for exactly $2g+1$ places $\omega$ that
(5) $\mathrm{val}_{\omega}(a_{k}-a_{\ell})=\begin{cases}1&\text{ if }\{k,\ell\}=\{i,j\},\\ 0&\text{ otherwise},\end{cases}$
and for all places $\omega$, and all $a_{i}$, we have $\mathrm{val}_{w}(a_{i})=0$ and $\mathcal{O}_{K}/(\omega)\geq B$.
We focus on understanding the Mordell-Weil rank of the family of quadratic twists of hyperelliptic curves that are BB-über-generic. Before doing so, we recall the following condition on $C$ from [BM, Condition 5.1] and [MS, Proposition 4.3]. Once again, we would like to thank Adam Morgan for suggesting the importance of the third condition provided below.
Condition 3.5 (Condition 5.1 from [BM] and Proposition 4.3 from [MS]).
Let $C$ be a hyperelliptic curve over $K$ with $K(\mathrm{Jac}(C)[2])=K$. Then $C$ satisfies the following three conditions.
  • The invariant subspaces of $\mathrm{Jac}(C)[2]$ with respect to $\mathrm{Gal}(K(\mathrm{Jac}(C)[4])/K)$ are 0 or $\mathrm{Jac}(C)[2]$.
  • The set of elements of $\mathrm{End}$ which commute with every element of $\mathrm{Gal}(K(\mathrm{Jac}(C)[4])/K)$ via multiplication is isomorphic to $\mathbb{F}_{2}$.
  • There exists a place $\omega$ of $K$ such that $C$ has a single nodal singularity at $\omega$. (In particular, the latter condition is satisfied if there exists a place $\omega$ and a tuple $(i,j)$ such that $\mathrm{val}_{\omega}(a_{i}-a_{j})=1$ and $\mathrm{val}_{\omega}(a_{k}-a_{l})=0$ for any other $k,l$ such that $\{k,l\}\neq\{i,j\}$.)
Remark 3.6.
We note that if $C$ is $B$-generic, then there exists a square-free element $d\in\mathcal{O}_{K}^{\times}$ such that $C_{d}$ satisfies Condition 3.5. To see this, we note that [BM, Proposition 8.10] shows that any hyperelliptic curve $C:y^{2}=(x-a_{1})\cdots(x-a_{2g+1})$ satisfying the following two conditions
  • There exists a set of places $\{\omega_{1},\omega_{2},\cdots,\omega_{2g}\}$ such that for each $1\leq i\leq 2g$, $\mathrm{val}_{\omega_{i}}(a_{i}-a_{2g+1})=1$ whereas $\mathrm{val}_{\omega_{i}}(a_{k}-a_{l})=0$ for any other $k,l$ such that $\{k,l\}\neq\{i,2g+1\}$.
  • There exists an element $\zeta\in\mathrm{Sel}_{2}(\mathrm{Jac}(C))\setminus\mathrm{Jac}(C)[2]$ such that the restriction of $\zeta$ over $H^{1}(K_{\omega_{i}},\mathrm{Jac}(C)[2])$ is unramified for all $1\leq i\leq 2g$.
satisfies the first two requirements in Condition 3.5. By [KM, Theorem 3.8], there exists a square-free element $d\in\mathcal{O}_{K}^{\times}$ such that the quadratic twist $C_{d}$ of a $B$-generic hyperelliptic curve $C$ satisfies the two bullet points mentioned above. By the condition on $a_{i}$’s, one can easily check that $C$ has a single nodal singularity at the place $\omega_{1}$, which satisfies the last condition from Condition 3.5.
Remark 3.7.
Let $C$ be a hyperelliptic curve over $\mathbb{Q}$. One can check the first two conditions from Condition 3.5 by computing the Weil pairing $e_{2}:\mathrm{Jac}(C)[2]\times\mathrm{Jac}(C)[2]\to\mu_{2}$. Suppose $\langle P_{1},P_{2},\cdots,P_{2g}\rangle$ is a basis for the set $\mathrm{Jac}(C)[2]$. If for every pair $1\leq i<j\leq 2g$ we have $e_{2}(P_{i},P_{j})=-1$, then the first two conditions hold by [BM, Proposition 2.15].
For the special case when $C$ is given by the Weierstrass model $C:y^{2}=(x-b_{1})\cdots(x-b_{4k+3})$ for some positive integer $k$, we can combine [Yu16, Corollary 7.6] and [Smi22b, Theorem 2.14] to ensure that at least $50\%$ of Jacobians of quadratic twists of $C$ have rank 0. In the case when $C$ is not of the form above, we can replace the third condition in Condition 3.5 with the requirement that there exists a single square-free element $d\in\mathbb{Q}^{\times}$ such that $\dim_{\mathbb{F}_{2}}\mathrm{Sel}_{2}(\mathrm{Jac}(C_{d})/\mathbb{Q})\equiv 0\mod 2$. This condition ensures that there is a positive density of Jacobians of rank 0, as shown in the discussion following after [Smi22b, Case 2.13] and the main theorem of [BM].
@@ 3.1. Higher genus - over any number field
Let $C:y^{2}=(x-a_{1})\cdots(x-a_{2g+1})$ be a hyperelliptic curve over a number field $K$ that satisfies Condition 3.5. Then there are infinitely many square-free elements $d\in\mathcal{O}_{K}^{\times}$ such that $\mathrm{Rank}(\mathrm{Jac}(C_{d})(K))=0$.
@@
The proof of [KM, Corollary 6.15] and the first paragraph of the proof of [MS, Proposition 4.3] implies that all the conditions in [Smi22b, Case 2.13] are satisfied, and that a positive density of square-free elements $d\in\mathcal{O}_{K}$ satisfies $\dim_{\mathbb{F}_{2}}\mathrm{Sel}_{2}(\mathrm{Jac}(C_{d})/K)\equiv 0\mod 2$. Hence, we can apply [Smi22b, Theorem 2.14] to show that there are infinitely many quadratic twists of $C$ such that $\mathrm{Rank}(\mathrm{Jac}(C_{d})/K)=0$. ∎
@@
For any $g\geq 2$, there are infinitely many hyperelliptic curves $C$ of genus $g$ over a number field $K$ such that $\mathrm{Rank}(\mathrm{Jac}(C)(K))=0$.